A mixture of 75 mol% methanol and 25 mol% methyl acetate is fed to a flash evaporator. The feed stream (liquid) is at 20 °C The flash evaporator runs at 55 °C, 90 kPa. What is the energy input required to operate this flash evaporator? ln(

0 0
Answer #1

gnen hot 2 X2 P68 813 KYa 2 A i AX 2 2 SScanned with CamScannerx1= [0 0.1 0.2 0.3 0.4 0.5 x2- 1-xl; T= 55+273; A=2.771-0.00523* T; gamma1= exp (A. * x2·2 ) ; gamma 2= exp (A.*x1.*2); Pisa110 105 - 100 95 90 85 80 75 70 65 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9x1-0.8; x2= 1-x1 ; T= 55+273; A=2 . 771-0. 00523* T; gamma l= exp (Ax22); gamma2= exp (A* x12); Plsat-68.813 P2 sat= 94 .> ch >> ch >>sd > sd yl 0.6072 0.3928 >>ch > sd y1 = 0.6072 y2 = 0.392829-g4 kJ / one make of r eth, and SScanned with CamScanner